# PIKAN: Physics-Inspired Kolmogorov–Arnold Networks for Explainable UAV Channel Modelling

### v1

@@ Abstract
Unmanned aerial vehicle (UAV) communications demand accurate yet interpretable air-to-ground (A2G) channel models that can adapt to non-stationary propagation environments. While deterministic models offer interpretability and deep learning (DL) models provide accuracy, both approaches suffer from either rigidity or a lack of explainability. To bridge this gap, we propose the Physics-Inspired Kolmogorov–Arnold Network (PIKAN) that embeds physical principles (e.g., free-space path loss, two-ray reflections) into the learning process. Unlike physics-informed neural networks (PINNs), PIKAN is more flexible for applying physical information because it introduces them as flexible inductive biases. Thus, it enables a more flexible training process. Experiments on UAV A2G measurement data show that PIKAN achieves comparable accuracy to DL models while providing symbolic and explainable expressions aligned with propagation laws. Remarkably, PIKAN achieves this performance with only 232 parameters, making it up to 37 times lighter than multilayer perceptron (MLP) baselines with thousands of parameters, without sacrificing correlation with measurements and also providing symbolic expressions. These results highlight PIKAN as an efficient, interpretable, and scalable solution for UAV channel modelling in beyond-5G and 6G networks.
@@ 1.1 Related Works
In parallel, deep learning (DL) approaches have emerged as alternatives for channel modelling. Architectures such as generative adversarial networks (GANs) have been proposed to generate realistic wireless channels and enrich training datasets [9]. Beyond generative models, physics-informed neural networks (PINNs) [10] have been introduced to embed physical equations directly into the training process. Therefore, they constrain the learning task with physical laws. PINNs promote consistency with known physical models; however, they can be overly rigid, struggling to adapt when empirical data deviates from theoretical assumptions. Other DL-based methods [11, 12, 13] have demonstrated strong performance but remain black-box in nature, offering no or very limited interpretability for underlying propagation mechanisms.
@@ 1.3 Contributions
We provide a path loss model considering other environmental parameters along with symbolic expressions since employing DL enables to learn the relation between path loss and other parameters than distance and carrier frequency, which are not considered in most of the deterministic models.
@@
We prefer to call it physics-inspired to differentiate it from PINN as our approach guides but does not rigidly constrain the learning process, while PINNs [10], which embed physical model in the loss function. Our approach provides more flexibility to capture real-world deviations from theory.
@@ 3 The Dataset
The dataset used in this study is collected with a custom-built software-defined radio (SDR) channel sounder integrated into the Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) testbed [15]. The measurements take place in a rural environment with minimal interference and few obstacles, which allows isolation of the key A2G channel characteristics.
A total of nine UAV flights are carried out using a hexacopter equipped with a portable receiver node that includes a USRP B210, a GNSS-disciplined oscillator (GNSS-DO) board, and an Intel NUC for control and data collection. A fixed transmitter node, mounted on a 10 m tower and equipped with a USRP B210 and GNSS-DO board, provides the reference signal. Both ends operate at a 56 MHz sampling rate, which is the maximum bandwidth supported by the USRP B210.
@@
where is the scaled feature, is the training-data matrix, denotes its -th feature column, is the median, and is the interquartile range (third minus first quartile) computed from the training data only. Using robust scaler when standardizing the features avoids from the impacts of outliers. The scaling parameters are computed exclusively from the training data to avoid information leakage. These parameters are then used to transform both the training and test datasets. The scaler parameters are given in Table 1.
@@ 5 Experiments
As noted in Section 4, we perform a grid search on hyperparameters and input features such as carrier frequency, distance, and so on to identify the best model–data combination111A full detailed discussion on the available features is provided in Appendix 8.. Sweeping and grid size , the lowest test RMSE is achieved at and . The most influential input features are the center frequency , horizontal distance , vertical distance , and angles of arrival and . Complete grid-search results are given in Table LABEL:tab:model_inputs_sweep. As shown in Fig. 6(a), as the grid , increases, the training performance generally improves while the test performance decreases. This is because the higher leads to a finer fit on the train data. Increasing behaves as a regularization because it leads to smoother splines. Our model selection is based on the balance between train performance and avoiding overfitting.
@@ 5.2 Physics-Inspired KANs
After the pruning, we fix certain activation functions to predetermined symbolic functions, considering the free space path loss model as given below:
@@
Besides FSPL, we utilize two-ray model [21].
@@
This inspiration reduces the MAE from to , which is lower than what both the base and the PIKAN-FSPL models achieve. Also, it provides the lowest RMSE and the best correlation with the measurements. Similar to PIKAN-FSPL, fixing the remaining activation functions to symbolics in the library creates a fully symbolic PIKAN-2R model as described in (18) with an MAE of . However, retraining the affine parameters leads to a higher error, which is remaining as an open issue for our next studies.
@@ 5.3.1 Deterministic Models
Table 3 presents the evaluation of deterministic baselines alongside the proposed symbolic KAN and PIKAN models. Classical deterministic models such as UMa LoS and UMa NLoS show limited accuracy, with large error magnitudes and correlation coefficients that are either low ( for UMa LoS) or unstable despite reduced RMSE. Other empirical models, such as FE2R and ITU-S, provide moderate improvements, but their errors remain an order of magnitude larger than the proposed methods.
@@ 6 Conclusions and Future Work
This work introduces physics-inspired Kolmogorov-Arnold network (PIKAN) for unmanned aerial vehicle (UAV) channel modeling, which combines the symbolic interpretability of KANs with the inductive guidance of physics-based models. Benchmarking against both deterministic baselines and deep MLPs shows that PIKAN provides the best balance between accuracy, complexity, and interpretability. For example, it achieves mean absolute error (MAE) 3.07 and with 37 fewer parameters than the best-performing multilayer perceptron (MLP). By providing symbolic path loss expressions that conform to physical intuition, PIKAN provides a new middle ground between physics-based models and black-box deep learning (DL) models.
@@ Acknowledgements.
The authors would like to thank Prof. İsmail Güvenç and Dr. Anıl Gürses for providing access to the UAV A2G measurement dataset used in this study. This work is supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada Alliance grant ALLRP 579869-22 ("Artificial Intelligence Enabled Harmonious Wireless Coexistence for 3D Networks (3D- HARMONY)").References

### v2

@@ Abstract
Unmanned aerial vehicle (UAV) communications demand accurate yet interpretable air-to-ground (A2G) channel models that can adapt to non-stationary propagation environments. While deterministic models offer interpretability and deep learning (DL) models provide accuracy, both approaches suffer from either rigidity or a lack of explainability. To bridge this gap, we propose the Physics-Inspired Kolmogorov–Arnold Network (PIKAN) that embeds physical principles (e.g., free-space path loss, two-ray reflections) into the learning process. Unlike physics-informed neural networks (PINNs), PIKAN is more flexible for applying physical information because it introduces them as adaptable inductive biases. Thus, it enables a more flexible training process. Experiments on UAV A2G measurement data show that PIKAN achieves comparable accuracy to DL models while providing symbolic and explainable expressions aligned with propagation laws. Remarkably, PIKAN achieves this performance with only 232 parameters, making it up to 37 times lighter than multilayer perceptron (MLP) baselines with thousands of parameters, without sacrificing correlation with measurements and also providing symbolic expressions. These results highlight PIKAN as an efficient, interpretable, and scalable solution for UAV channel modelling in beyond-5G and 6G networks.
@@ 1.1 Related Works
In parallel, deep learning (DL) approaches have emerged as alternatives for channel modelling. Architectures such as generative adversarial networks (GANs) have been proposed to generate realistic wireless channels and enrich training datasets [9]. Beyond generative models, physics-informed neural networks (PINNs) [10] have been introduced to embed physical equations directly into the training process. PINNs promote consistency with known physical models; however, they can be overly rigid, struggling to adapt when empirical data deviates from theoretical assumptions. Other DL-based methods [11, 12, 13] have demonstrated strong performance but remain black-boxes in nature, offering no or very limited interpretability for underlying propagation mechanisms.
@@ 1.3 Contributions
We provide a path loss model considering other environmental parameters along with symbolic expressions because employing DL enables to learn the relation between path loss and other parameters than the distance and carrier frequency, which are not considered in most of the deterministic models.
@@
Unlike PINNs [10], where physical phenomena appear directly in the loss function, our approach employs no such regularizer. Instead, we configure the KAN by fixing its activation functions to symbolic relations from physical expressions, such as the free space path loss (FSPL). Thus, we refer to the model as inspired to distinguish it from PINNs.
@@ 3 The Dataset
The dataset used in this study is collected with a custom-built software-defined radio (SDR) channel sounder integrated into the Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) testbed [15, 19]. The measurements take place in a rural environment with minimal interference and few obstacles, which allows isolation of the key A2G channel characteristics.
A total of nine UAV flights are carried out using a hexacopter equipped with a portable receiver node that includes a USRP B210, a GNSS-disciplined oscillator (GNSS-DO) board, and an Intel NUC for control and data collection. A fixed transmitter node, mounted on a 10 m tower and equipped with a USRP B210 and a GNSS-DO board, provides the reference signal. Both ends operate at a 56 MHz sampling rate, which is the maximum bandwidth supported by the USRP B210.
@@
where is the scaled feature, is the training-data matrix, denotes its -th feature column, is the median, and is the interquartile range (third minus first quartile) computed from the training data only. Using a robust scaler when standardizing the features avoids from the impacts of outliers. The scaling parameters are computed exclusively from the training data to avoid information leakage. These parameters are then used to transform both the training and test datasets. The scaler parameters are given in Table 1.
@@ 5 Experiments
As noted in Section 4, we perform a grid search on hyperparameters and input features such as carrier frequency, distance, and so on to identify the best model–data combination. A full detailed discussion on the available features is provided in Appendix 8. Sweeping and grid size , the lowest test RMSE is achieved at and . The most influential input features are the center frequency , horizontal distance , vertical distance , and angles of arrival and . Complete grid-search results are given in Table LABEL:tab:model_inputs_sweep. As shown in Fig. 6(a), as the grid , increases, the training performance generally improves while the test performance decreases. This is because the higher leads to a finer fit on the train data. Increasing behaves as a regularization because it leads to smoother splines. Our model selection is based on the balance between train performance and avoiding overfitting.
@@ 5.2 Physics-Inspired KANs
After the pruning, we fix certain activation functions to predetermined symbolic functions, considering the FSPL model as given below:
@@
Besides FSPL, we utilize two-ray model [22]:
@@
This insight reduces the MAE from to , which is lower than what both the base and the PIKAN-FSPL models achieve. Also, it provides the lowest RMSE and the best correlation with the measurements. Similar to PIKAN-FSPL, fixing the remaining activation functions to symbolics in the library creates a fully symbolic PIKAN-2R model as described in (18) with an MAE of . However, retraining the affine parameters leads to a higher error, which remains an open issue for our next studies.
@@ 5.3.1 Deterministic Models
Table 3 presents the evaluation of deterministic baselines alongside the proposed symbolic KAN and PIKAN models. Classical deterministic models such as UMa LoS and UMa NLoS show limited accuracy, with large error magnitudes and correlation coefficients that are either low ( for UMa LoS) or unstable despite reduced RMSE. Other empirical models, such as FE2R and ITU-S, provide moderate improvements, but their errors remain an order of magnitude larger than those of the proposed methods.
@@ 6 Conclusions and Future Work
This work introduces physics-inspired Kolmogorov-Arnold network (PIKAN) for unmanned aerial vehicle (UAV) channel modelling, which combines the symbolic interpretability of KANs with the inductive guidance of physics-based models. Benchmarking against both deterministic baselines and deep multilayer perceptrons (MLPs) shows that PIKAN provides the best balance between accuracy, complexity, and interpretability. For example, it achieves mean absolute error (MAE) 3.07 and correlation, with 37 fewer parameters than the best-performing MLP. By providing symbolic path loss expressions that conform to physical intuition, PIKAN provides a new middle ground between physics-based models and black-box deep learning (DL) models.
@@ Acknowledgements.
The authors would like to thank Prof. İsmail Güvenç and Dr. Anıl Gürses for providing access to the UAV A2G measurement dataset used in this study. This work is supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada Alliance grant ALLRP 579869-22 (”Artificial Intelligence Enabled Harmonious Wireless Coexistence for 3D Networks (3D- HARMONY)”).References
@@
- [19] A. H. F. Raouf, D. Lee, M. Rahman, S. Masrur, G. Reddy, C. Dickerson, M. S. Hossen, S. V. Villar, A. Gürses, S. Singh et al., “Wireless datasets for aerial networks,” arXiv preprint arXiv:2510.08752, 2025.